# Resource-constrained Fairness

### v4

@@ 2.3 Cost of Fairness
Before we proceed further, we should note that assigning a cost to actions is inherently a political action, and that often these costs reflect the beliefs of data scientists and other professionals as much as they do the raw data. This is particularly apparent in the case of personal loans.
While banks have good knowledge of the repayment habits of people they have lent money to in the past, they have much less knowledge of how people they did not lend money to might repay. In no small part, this knowledge asymmetry has allowed for the persistence of institutional redlining whereby loans are routinely denied to people in black or Hispanic dominant neighbourhoods when they would be granted to individuals in similar circumstances that live in white dominant neighbourhoods  (Markup, 2021). Similarly, while it is tempting to assume that medical data is reliable, and for example, that we are forecasting the result of a future biopsy that will infallibly determine if cancer is present or not, for many problems the information is less clear-cut, and ground-truth may come from two experts attempting to diagnose from limited post-mortem records, with a third expert adjudicating in the case of disagreement Lee et al. (2023).
Indeed, one justification for constraints such as demographic parity is that ground-truth data is often gender- or ethnically-biased, and in some circumstances we can a priori expect uniform rates across these populations (in line with the “We’re All Equal” world-view of Friedler et al. (2021), which asserts that there are no innate differences between groups). As such, any estimate of cost comes with the usual caveat of assuming the ground-truth is correct. But even as simplified approximations, these costs remain useful for understanding the potential trade-offs in enforcing fairness.
With this in mind, we consider the following question: What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
Under constrained resources, this leads to the same results as enforcing fairness, i.e. equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
@@
While other studies have previously analyzed the cost of fairness (Friedler et al., 2019; Haas, 2019; von Zahn et al., 2021), they did not consider the implicit trade-off that comes from constrained resources in decision-making, but instead measured how classifiers deteriorate with fairness.
@@
We might also be interested in performance metrics outside the fairness metric enforced. To quantify the loss in precision, we ask: What is the change in precision if we minimize the harm of the worst-off group? This represents a common scenario. In a medical context it corresponds to the question: What proportion of healthy patients instead of sick will we see if we increase test sensitivity for the worst off groups? In the results, we measure this cost and its contributory factors. This leads to a follow-up question: What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we minimise the harm of the worst off-group? The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increases the resources available, and target these resources at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources. The number of available resources can be an alterable parameter that should explicitly be taken into account.
3 Resource-constrained Fairness and its Cost
@@
While standard fairness methods do not enforce fairness as the selection rate varies (see Figure 2), there is a straightforward algorithm for enforcing fairness at any selection rate, based on the harms based analysis of section 2.2. Given an existing classifier $c_{w}$, we say that a classifier makes a positive decision regarding datapoint $x$ if $c_{w}(x)>t$, where t is some threshold. We consider fairness as in Equation LABEL:eq:fair_class where some notion of harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$is being equalized across groups.
Now, we simply vary a per group threshold $t_{g}$, while computing the group harm $\mathrm{H}_{g}[c_{w}]$, and the corresponding selection rate $\mathrm{r}_{g}[c_{w}]$. We consider some $h$, and select the set of thresholds so that the harm per group lies just under this $h$. This is (approximately) fair with respect to harm $H$, and has a global selection rate
$r=\frac{\sum_{g}|g|\,\mathrm{r}_{g}[c_{w}>t_{g}]}{\sum_{g}|g|}$ (6)
where $|g|$ is the number of datapoints in group $g$. By sweeping over possible values $h$, and examining the corresponding global selection rate, we can find a fair solution that lies just under a target selection rate. The corresponding thresholds can be used to enforce fairness on previously unseen test data.
@@
The resource allocation process is particularly straightforward for demographic parity. In this case the harm corresponds to 1- the selection rate, and the process simplifies to just selecting the top $r$ from each group.  [^4] Note that in particular case of demographic parity ideal thresholds can be calculated on the test set, but for most other measures the thresholds need to be calculated on a separate validation set, as they require access to the target label in order to compute the harms. Owing to sampling error, thresholds selected on the validation data do not perfectly correspond to selection rates on unseen data, and to ensure that the number of instances predicted as positive exactly matches the available resource level, we do not directly use the thresholds. Instead, we calculate the proportion of resources to allocate to each group for each resource level $R$ on the validation set, and then enforce this on the test set.
Our strategy belongs to the class of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained (Mehrabi et al., 2021). We conduct this analysis across a complete range of selection rates $r$ (from 1% to 100 %) to represent different resource levels.
Given this approach, we can now quantify the cost of fairness, defined as the difference in performance (measured as either precision or recall) between the default and fair labeling approaches, for a fixed number of selected instances (R). We vary the available resources $R$ (and consequently the selection rate $r$) and investigate how this influences the cost of fairness.
3.1 Datasets and Classifiers
We use several real-world datasets common in tabular fair machine learning (Le Quy et al., 2022). We also extend our findings to two datasets (CelebA Liu et al. (2015) and Fitzpatrick17k Groh et al. (2021)) from computer vision. The Adult Income dataset is taken from from the 1994 census data, and contains a prediction task identifying if an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. (2020) and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We prepossess the data and binarize the labels following the approach of Zong et al. (2022). We specify the protected and target attributes in Table 1. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $ns$ representing the protected group and $s$ the advantaged group it can be mathematically expressed as:
$P(Y=1\mid S=ns)-P(Y=1\mid S=s)$
[…]
Classifiers
For the tabular datasets, we use eXtreme Gradient Boosting (XGBoost) Chen and Guestrin (2016). We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. (2016) pretrained on ImageNet Deng et al. (2009) for feature extraction.
3.2 Metrics
3.2.1 Performance metrics
$P(S(x_{i})>S(x_{z})\mid y_{i}=1,y_{z}=0)$
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall and accuracy at $R$ for a specific dataset and model and a fixed value of $R$ (Rodolfa et al., 2021). We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall [^5] The proportion of actual positives in the top $R$ over the total number of actual positives in Table 2 and Figure 5.
3.2.2 Fairness Metrics
We focus on two widely recognized fairness measures used to assess disparities between groups.
Demographic parity (also known as statistical parity) requires that the rate of positive decisions is roughly equal for both the protected group and the privileged group:
$P(\hat{Y}=1|S=s)\approx P(\hat{Y}=1|S=ns)$
Equal opportunity requires the true positive rate to be approximately the same across groups (Hardt et al., 2016), which enforces equal recall:
$P(\hat{Y}=1|S=s,Y=1)\approx P(\hat{Y}=1|S=ns,Y=1)$
4 What influences the Cost of Fairness?
Table 2: Model performance (entire population, privileged group and protected group) and the average cost of fairness (the loss in precision and recall) when enforcing DP and EO. The average is calculated over all the selection rates from $1\%$ to $100\%$.
Dataset Adult Compas Dutch Law CelebA Fitzpatrick17K
AUC 0.899 0.811 0.917 0.871 0.957 0.826
$AUC_{priv}$ 0.884 0.797 0.884 0.852 0.931 0.829
$AUC_{prot}$ 0.887 0.793 0.914 0.842 0.969 0.814
Avg. loss in precision (DP) 0.024 0.025 0.028 0.006 0.069 0.001
Avg. loss in precision (EO) 0.010 0.013 0.007 0.004 0.009 0.000
Avg. loss in recall (DP) 0.024 0.016 0.028 0.004 0.062 0.001
Avg. loss in recall (EO) 0.012 0.007 0.008 0.003 0.013 0.001
@@ 4 What influences the Cost of Fairness?
In Table 2, we analyze the AUC performance of our machine learning model for different groups: the entire population, and separately for protected and privileged groups. Additionally, we report the average cost of enforcing fairness as the loss in both precision and recall between the default allocation and the fair allocation, averaged over all selection rates. We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity. The average cost is also very low for the Law dataset, despite its significant base rate disparity. A possible reason for this could be the relatively small size of the protected group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset.
Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the protected group, the noisiness of the whole dataset and the noisiness of the protected group. The impacts of these changes are visualized in Figure 2, with the original dataset’s results represented by a black line in each figure.
@@
Figure 2: Impact of parameters on the cost of fairness, measured as the loss in precision (Adult Income dataset). We see that reducing the base rate disparity leads to a lower cost of fairness. Similarly, introducing more noise across all groups, also lowers the cost of fairness, but introducing more noise exclusively in the protected group, increases the cost of fairness. Finally, reducing the size of the protected group, results in a decrease in the cost of fairness as well.

### v5

@@ 3.1 Cost of Fairness
Before we proceed further, we note that assigning a cost to actions is inherently a political action, and that often these costs reflect the beliefs of data scientists and other professionals as much as they do the raw data.
Indeed, one justification for constraints such as demographic parity is that ground-truth data is often gender- or ethnically-biased, and in some circumstances we can a priori expect uniform rates across these populations (in line with the “We’re All Equal” world-view of Friedler et al. (2021), which asserts that there are no innate differences between groups). As such, any estimate of cost comes with the usual caveat of assuming the ground-truth is correct. But even as simplified approximations, these costs remain useful for understanding the potential trade-offs in enforcing fairness. With this in mind, we consider the following question: What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
Under constrained resources, this has the same result as enforcing equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
@@
While other studies have previously analyzed the cost of fairness Corbett-Davies et al. (2017); Friedler et al. (2019); Haas (2019); von Zahn et al. (2021), they did not consider the implicit trade-off that comes from constrained resources in decision-making, but instead measured how classifiers deteriorate with fairness. Some of the calculated costs may very well arise from selecting a different number of people; however, we will keep this number constant and focus instead on the costs associated with selecting a different set of individuals.
@@
We might also be interested in performance metrics outside the fairness metric enforced. Of particular interest is the change in precision that occurs if we minimize the harm to the worst-off group. This represents a common scenario. In a medical context, it corresponds to the question: What proportion of healthy patients instead of sick will we see as we increase test sensitivity for disadvantaged groups?
This leads to a follow-up question: What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we enforce equality? The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increase the resources, and target them at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources.
4 Bounding the cost of fairness
4.1 Bounding the Cost of Fairness
We now formalize these costs and consider the cost of the change induced by enforcing fairness at a particular selection rate. As we are modeling the cost of inducing fairness under rate constraints, we consider the changes in the labeling induced by swaps, where the label assigned to one instance changes from positive to negative, and the label assigned to another instance changes from negative to positive. A sequence of swaps must always preserve the global selection rate, and any pair of labellings with the same selection rate, can be transformed from one to another by a sequence of swaps. We write $p$ for the proportion of instances in the dataset swapped from positive to negative (or equivalently the proportion swapped from negative to positive), and $c$ for the average cost of swapping a pair.
By definition, the cost of inducing fairness is:
$\text{cost}=p\cdot c$ (7)
Note that $p\in[0,0.5]$, and for standard linear measures (e.g. accuracy, recall, specificity) the second term is bounded by
$c=\frac{\text{\#instances in dataset}}{\text{\#instances used in the measure}}$ (8)
i.e. $c=1$ for accuracy, and, writing $b$ for the label base rate, $c=b^{-1}$ for Recall and False Positive Rate, and $c=(1-b)^{-1}$ for Specificity and False Negative Rate. As such, $pc$ and $c/2$ are both upper bounds on the cost of fairness.
For example, at a given global selection rate $r$, the cost of fairness, in terms of change in accuracy, is at most the proportion of swaps ($p$); while the cost in terms of change of recall is at most $p/b$.
We can create successive bounds based on this identity: Given $r$, $p$ is bounded above by $r$ (all positive instances are swapped); and $(1-r)$ (all negative instances are swapped). Therefore, the cost of fairness is bounded above by $rc$, and $(1-r)c$. As such, if $r$ or $(1-r)$ is much smaller than $c^{-1}$, there is little cost in enforcing fairness.
This matches our empirical finding in Section 5 that the cost is typically bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness Menon and Williamson (2018); Friedler et al. (2019); Hort et al. (2023) overestimate the cost of fairness for scenarios with other resource levels.
For scenarios where the selection rate substantially exceeds the base rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to. These bounds are much weaker for non-linear measures such as precision, if they exist. For precision, assuming a fixed global selection rate $r$, we have $c=r^{-}1$. Consequentially, the low selection-rate bound given by $rc=1$ is vacuous. However, other bounds remain informative when there is a higher selection rate – particularly, the bounds $(1-r)c^{-1}$ and $gc^{-1}$, indicate that we should see similar patterns in the costs with respect to precision for higher selection rates. Empirically, we find this — precision exhibits similar trends to the linear measures of recall and accuracy. Reflecting the much weaker bound for low selection rates, it sometimes exhibits strong changes at low resource levels, which are not present for accuracy and recall. This can be seen in Figures 3(b) and 3(d).
Similarly, given $g$, the proportion of the dataset occupied by the smallest group, we know $p<g$, and the cost of inducing fairness is bounded above by $gc$ [^2] Relabelling only a subset of the smallest group is not always sufficient to enforce fairness, e.g. consider equal precision, but it holds for linear measures. Again, if $g$ is much smaller than $c$, there is little cost in inducing fairness. This finding is supported in our experiments (see Figures 2(g) and 2(h)).
Finally, looking at the second term $c$ (the average cost of relabeling), we find that this second factor explains much of what we will see in Figure 2. For example, increasing global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the average cost of relabelling points. Similarly, increasing classification noise for the group with greatest classification error, or increasing the difference in base rates, increases the performance gap between groups and consequentially the cost of relabelling.
Following this analysis, it is reasonable to ask if bounds exist for levelling up, and if there are bounded increases in selection rate for a disadvantaged group guaranteed to remove unfairness (while keeping the selection rate of the advantaged group as-is). In brief, the answer is yes for demographic parity; for equal opportunity you may need to relabel all datapoints in a disadvantaged group; and levelling up need not be possible for equal precision.  [^3] These results differ from Section 4.1 because the previous analysis was advantaged by the stability of measures – we altered per group selection rates and hoped that global measures remained stable. In this new case, we are trying to alter per-group measures and the stability of measures is now a disadvantage.
For demographic parity, there are no stability concerns. Let $g$ be the proportion of the dataset corresponding to the disadvantaged group, and $r$ the global classifier selection rate. Then an upper bound for the selection rate excluding members of the disadvantaged group is $r/(1-g)$, and at most we require an additional $gr/(1-g)$ proportion of points to be positively labelled to obtain parity. For equal opportunity, only instances assigned a positive ground-truth label alter the recall, and it is possible that a badly-trained classifier selects all negatively labeled instances first, and only then positively labelled points. If we consider a classifier like this, and a group with only one positively labelled point (in the ground truth), all points within the group must be selected to get any non-zero recall rate for that group. As such, $g$ is a bound. For equal precision, we consider the same example, and note that while you must select all datapoints to obtain non-zero precision, the precision will be the base rate of the group, and it may be insufficient for equality with respect to the other group.
4.2 Varying Global Selection Rates
Until now, we assumed the resources to be fixed, but practitioners may find it feasible to increase or decrease them, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Donahue and Kleinberg (2020) discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role.
5 Enforcing Fairness
Fairness at different selection rates can be enforced using two existing approaches. Kwegyir-Aggrey et al. (2023) enforces equal opportunity and demographic parity by aligning per-group score distributions while minimizing the earthmovers difference. The fairness toolkit of Delaney et al. (2024) optimizes an objective subject to a constraint (such as maximize min group recall, subject to the overall selection rate being less than 0.7). All methods fall into the family of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained Mehrabi et al. (2021). For completeness, we set out our strategy for efficiently finding thresholds, however, this has limited novelty and should not be considered a significant contribution.
@@ 5 Enforcing Fairness
To measure the cost of fairness, we implement a per-group parameter sweep, based on the harms-based analysis of Section 2. Given a classifier $c_{w}$, we say it makes a positive decision regarding datapoint $x$ if $c_{w}(x)>t$, where $t$ is some threshold. We consider fairness as in Equation LABEL:eq:fair_class where some notion of harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$  is being equalized across groups. Now, we simply vary a per group threshold $t_{g}$, while computing the group harm $\mathrm{H}_{g}[c_{w}]$, and the corresponding selection rate $\mathrm{r}_{g}[c_{w}]$. We consider some measure of acceptable harm $h$, and select the set of per group thresholds $t_{g}$ so that the harm per group lies just under this $h$. This is (approximately) fair with respect to harm $H$, and has a global selection rate
$r=\frac{\sum_{g}|g|\cdot\mathrm{r}_{g}[c_{w}>t_{g}]}{\sum_{g}|g|}$ (9)
@@
where $|g|$ is the number of datapoints in group $g$. By sweeping over possible values $h$, and examining the corresponding global selection rate, we find a fair solution that lies just under a target selection rate. The corresponding thresholds can be used to enforce fairness on previously unseen test data. We conduct an analysis across a range of selection rates $r$ (from 1% to 100 %) representing different resource levels. This process is particularly straightforward for demographic parity. In this case, the harm corresponds to 1- the selection rate, and the process simplifies to just selecting the top $r$ from each group. [^4] For demographic parity, ideal thresholds can be calculated on the test set, but for most other measures the thresholds need to be calculated on a separate validation set, as they require access to the target label in order to compute the harms. Owing to sampling error, thresholds selected on the validation data do not correspond perfectly to selection rates on unseen data. As such, to ensure that the number of instances predicted as positive exactly matches the available resources, we calculate the proportion of resources to allocate to each group for each resource level $R$ on the validation set, and transfer these proportions to the test set.
6 The Empirical Cost of Fairness
@@
We discuss the datasets, classifiers and metrics in Appendix A [^5] The code for the experiments is publicly available at https://github.com/SofieGoethals/RCF. In Table 1, we analyze the AUC performance of a machine learning model for different groups: the entire population, and separately for disadvantaged and advantaged groups. Additionally, we report the average cost of enforcing fairness as the loss in precision between the default allocation and the fair allocation, averaged over all selection rates. [^6] The results for recall and accuracy are in line with precision and can be found in Table 3 We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity (see Table 2). The average cost is also very low for the Law dataset, despite its significant base rate disparity. A possible reason for this could be the relatively small size of the disadvantaged group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset.
Table 1: Model performance (entire population, advantaged group and disadvantaged group) and the average cost of fairness (loss in precision) when enforcing DP and EO. The average is calculated over all the selection rates from $1\%$ to $100\%$.
Dataset Adult Compas Dutch Law CelebA Fitz17K AUC 0.896 0.812 0.917 0.870 0.957 0.826 $\text{AUC}_{priv}$ 0.887 0.797 0.884 0.847 0.931 0.829 $\text{AUC}_{prot}$ 0.871 0.793 0.914 0.858 0.969 0.814 Avg. cost (DP) 0.025 0.026 0.028 0.005 0.069 0.001 Avg. cost (EO) 0.009 0.014 0.007 0.003 0.009 0.000
@@
Figure 2: Impact of parameters on the cost of fairness, measured as the loss in precision (Adult Income dataset). We see that reducing the base rate disparity leads to a lower cost of fairness. Similarly, introducing more noise across all groups, also lowers the cost of fairness, but introducing more noise exclusively in the disadvantaged group, increases the cost of fairness. Finally, reducing the size of the disadvantaged group, results in a decrease in the cost of fairness as well.
Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the disadvantaged group, the noisiness of the whole dataset and the noisiness of the disadvantaged group. The impacts of these changes are shown in Figure 2, with the original dataset’s results represented by a black line.
Base rate disparity First, we alter the base rate disparity. We see in Figures 2(a) and 2(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target.
Noise We alter the performance of the model by introducing random noise (with varying degrees from 0 to 0.5) to the feature values of the whole dataset. The base rates of both groups remain the same, but the model struggles more with correctly classifying individuals, resulting in a lower AUC score. We see that as model performance decreases, the average cost of fairness also goes down, both for DP and EO (Figures 2(c)-2(d)). This global decrease shrinks the gap between the performances of each group, reducing the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 1. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^7] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost of enforcing demographic parity would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity, and is reached at a point between the base rates.
Subgroup noise We can also add noise only to the members of the disadvantaged group. This lowers the performance of the model for the disadvantaged group, but not for members of the advantaged group. As shown in Figures 2(e) and 2(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
Size of the disadvantaged group Lastly, the size of the disadvantaged group substantially influences the cost of fairness. When the disadvantaged group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 2(g) and 2(h), where reducing the size of the disadvantaged group, by randomly excluding part of this group, consistently lowers the cost of fairness. This is an expected finding and aligns with our bounds analysis in Section 4.1, however, we have not seen it discussed previously.
6.1 The Impact of Selection Rate